OpenStudy (anonymous):

general value of x which satisfy the equation 2sinx+1=0 ?

OpenStudy (anonymous):

sin x= 1/2

OpenStudy (anonymous):

You subtract one and then div by 2. So it should be sin x= -1/2

OpenStudy (anonymous):

so what would be general value

OpenStudy (anonymous):

\[\frac{ 7\pi }{ 6 } \pm 2\pi n\]

OpenStudy (anonymous):

\[\frac{ 11\pi }{ 6 } \pm 2\pi n\]

OpenStudy (anonymous):

Those would be your general solutions

OpenStudy (anonymous):

in this n is natural number but i don't have these in any of its options

OpenStudy (anonymous):

What are your options?

OpenStudy (anonymous):

Yes, n would be the natural number

OpenStudy (anonymous):



OpenStudy (anonymous):

options are a)n(pi)- (-1)^n(pi/4) b)n(pi)- (-1)^n(pi/6) c)n(pi)+ (-1)^n(pi/3) d)none for all n belongs to integers

OpenStudy (anonymous):

well with the -1/2. You need to look on the circle graph with all of the sin, cos, tan, and cot coordinates. If both of your general solutions are divided by 6 then I would say b. But the equations their I have never seen them that way before.

OpenStudy (anonymous):

i had marked the ans as b)

OpenStudy (anonymous):

Thats great! I believe it will be tahat

OpenStudy (anonymous):

I think it will be that one. It goes with my gerneal solutions that i found

OpenStudy (anonymous):

please reply i am closing the question so that i can ask more

OpenStudy (anonymous):

ooooooook

OpenStudy (anonymous):

i am confused in d)none or b)